OpenStudy (anonymous):

**i give medals** 3x+4y=11 x+3y=11 Use at least two points to graph each line. Label each point. Label each line.

OpenStudy (anonymous):

@tom982

OpenStudy (anonymous):

Let's start by solving for y in both equations: y = 11/4 - 3x/4 y = 11/3 - x/3

OpenStudy (anonymous):

After that we can solve for y(0) for both equations to find the y intersect. y = 11/4 so point 1 (0 , 11/4) y = 11/3 point 1 ( 0 , 11/3)

OpenStudy (anonymous):

we can also use this to find the x intersections for both lines x = 11/3 so point 2 ( 11/3, 0) x = 11 so point 2 ( 11, 0)

OpenStudy (anonymous):

There's your two points for each line to graph them, what else do you need?

OpenStudy (anonymous):

thx ill post the next 2 question and tag u

OpenStudy (anonymous):

Alright